# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v1

@@ Abstract.
In this paper, we analyze nonlinear model predictive control (MPC) using data-driven surrogates in the prediction and optimization step. First, we establish asymptotic stability of the origin, a controlled steady state, w.r.t. the MPC closed loop without stabilizing terminal conditions. To this end, we prove that cost controllability of the original system is preserved if proportional bounds on the approximation error hold. Here, proportional refers to state and control, while the respective constants depend on the approximation accuracy. The proportionality of the error bounds is a key element to derive asymptotic stability in presence of modeling errors and not only practical asymptotic stability. Second, we exemplarily verify the imposed assumptions for data-driven surrogates generated with kernel extended dynamic mode decomposition based on the Koopman operator. Hereby, we do not impose invariance assumptions on finite dictionaries, but rather derive all conditions under non-restrictive data requirements. Finally, we verify our findings with numerical simulations.
@@ 1. Introduction
Model predictive control (MPC) is an optimization-based control algorithm, that optimizes the future trajectory of the system on a finite horizon and defines a state feedback law by applying to the system the first element of the computed optimal input sequence RawlMayn17 . When e.g. a first-principles-based description of the system under control is not available, data-driven methods can be used to obtain a surrogate model for the prediction and optimization step. For linear systems, a popular data-driven approach BerbKohl20 is based on the use of Willems’ fundamental lemma WillRapi05 avoiding the model identification step and directly using input-output data markovsky2021behavioral ; FaulOu23 . For nonlinear systems, data-driven surrogate models can be obtained, e.g., by means of nonlinear ARX models denicolao1997stabilizing , neural networks schimperna2024robust-lstm ; ren2022tutorial , Bayesian identification piga2019performance and Gaussian process regression hewing2019cautious ; koller2018learning .
@@
A numerically and analytically appealing approach is Extended Dynamic Mode Decomposition (EDMD williams2015data ) based on the Koopman operator. The Koopman operator is a linear but infinite dimensional operator that encodes the behavior of the associated nonlinear dynamical system mezic2005spectral ; rowley2009spectral . Leveraging this linearity and for autonomous systems, EDMD provides a data-driven approximation of the compression of the Koopman operator to a finite-dimensional subspace spanned by the so-called observable functions Mezi22 . Various approaches extend this idea to control systems. In EDMD with control (EDMDc; proctor2016dynamic ; korda2018linear ), a surrogate control system that is linear in the (lifted) state and control is used. However, as shown in iacob:toth:schoukens:2022 , this method is often insufficient to capture direct state-control couplings. Bilinear models as in peitz:otto:rowley:2020 offer a reliable alternative that is underpinned by finite-data error bounds nuske2023finite . Besides ensuring a sufficient amount of data, a key challenge in EDMD is the choice of the observable functions. Here, data-driven approaches, such as neural networks yeung2019learning or kernel EDMD klus2020kernel provide a remedy. In kernel EDMD, the dictionary consists of the canonical features of an a-priori chosen kernel function centered at the data points. For this method, it is possible to build upon the solid approximation-theoretic foundation via reproducing kernel Hilbert spaces (RKHSs) to obtain finite-data error bounds kohne2024infty leveraging Koopman invariance of the underlying RKHS. This method and its error bounds have been extended to control systems in BoldPhil24 . An alternative approach, that shares the advantage of flexible state-control sampling, was proposed in bevanda2024nonparametric using an additional kernel function to express the dependency on the control input. So far, however, no error bounds are available for this method.
@@
The use of data-driven models gives rise to approximation errors, that are related to the choice of the model class used for approximation and the finite amount of data available for the identification. While the asymptotic stabilization of a steady-state equilibrium can be guaranteed in the nominal case, i.e. when the model used in the optimization step coincides with the system under control, in presence of modeling errors typically only practical asymptotic stability is achieved. Practical asymptotic stability means that the tracking error decreases until a certain threshold, and then stagnates to a value different from zero. The stagnation value can then usually be related proportionally to the underlying approximation error. Practical asymptotic stability of MPC with surrogate models based on EDMD has been shown both without and with the use of terminal ingredients in bold2024data and WortStra24 , respectively. In BoldScha25 , the results of bold2024data have been extended to kEDMD surrogate models. A method to obtain offset-free tracking in presence of modeling errors and asymptotically constant disturbances is so-called offset-free MPC pannocchia2003disturbance ; morari2012nonlinear . This technique has been successfully applied to MPC with Koopman operator-based surrogates. In particular, son2022development and chen2022offset consider linear offset-free MPC with surrogate models based on EDMDc, while SchiBold25 considers bilinear EDMD models and addresses the case in which a full description of the reference steady state is not available.
@@
The first contribution of this work is a general framework for stability analysis of MPC with data-driven surrogate models subject to approximation errors. Our first contribution is to show that if the error of the surrogate can be bounded in a proportional way, the MPC algorithm designed using the data-driven surrogate model asymptotically stabilizes the system. With proportional we mean that the error vanishes at the desired set point, and its magnitude is bounded by a quantity that depends on the size of the state and input. We show that proportional error bounds imply the preservation of cost controllability in data-driven approximations provided a sufficiently high approximation accuracy. Then, invoking this controllability property, we rigorously show asymptotic stability of the origin w.r.t. the MPC closed loop using the surrogate model in the prediction and optimization step, where the origin serves as a prototypical controlled steady state.
@@
The second main contribution of this paper is to verify the assumptions on the data-driven surrogate leveraging recently proposed finite-data error bounds on kernel EDMD under non-restrictive assumptions. To this end, we consider the learning framework proposed in BoldPhil24 , suitably modified to ensure that the model is consistent with the system dynamics in the origin. We derive proportional and uniform error bounds – without requiring invariance assumptions on finite dictionaries and under non-restrictive data requirements. Finally, our findings are verified with numerical simulations. [^1] We point out that this part is an extension of our accepted conference paper BoldScha25 now properly embedded into a more general learning framework such that we were able to infer asymptotic stability (instead of its only practical asymptotic stability).
@@
[…]
@@ 2. Data-driven MPC formulation
with state $x\in\mathbb{R}^{n}$ and input $u\in\mathbb{U}$, where $\mathbb{U}\subseteq\mathbb{R}^{m}$ is a convex and compact set containing the origin in its interior, i.e., $0=0_{\mathbb{R}^{m}}\in\operatorname{int}(\mathbb{U})$. The system dynamics $f$ is assumed to be continuous and locally Lipschitz continuous in its first argument, i.e., for each compact set $S\subset\mathbb{R}^{n}$ there is a Lipschitz constant $L_{f}=L_{f}(S)$ such that
$\|f(x,u)-f(y,u)\|\leq L_{f}\|x-y\|,\qquad\forall\,u\in\mathbb{U}.$ (2)
@@
Let the origin be a controlled equilibrium of the system (1), i.e. $f(0,0)=0$. The objective is to control the system to the origin with MPC taking the input constraints $u\in\mathbb{U}$ into account. However, since the system dynamics (1) is considered unknown, the MPC design is based on data-driven surrogate models
$x^{+}=f^{\varepsilon}(x,u),$ (3)
@@
with $\varepsilon\in(0,\bar{\varepsilon}]$, $\bar{\varepsilon}\in(0,\infty)$. The superscript $\varepsilon$ refers to the approximation accuracy and, throughout the paper, will be used to refer to all the parameters related to the surrogate dynamics (3). Typically, the approximation error depends on the available data (quantity, distribution etc.). To be more precise, we consider models accompanied by proportional error bounds, where proportional means that the error is zero at the origin and may grow proportionally with the size of state and input. This property is formalized in the following assumption.
Assumption 1 (Proportional error bound on a set $S$).
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior. Let the proportional error bound
$\|f(x,u)-f^{\varepsilon}(x,u)\|\leq c_{x}^{\varepsilon}\|x\|+c_{u}^{\varepsilon}\|u\|$ (4)
@@
hold for all $x\in S$, $u\in\mathbb{U}$, and $\varepsilon\in(0,\bar{\varepsilon}]$ for the approximation (3) of system (1), where the parameters $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$ satisfy $\lim_{\varepsilon\searrow 0}\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}=0$.
@@
Assumption 1 requires that the surrogate model is exact at the origin, which is a reasonable assumption since the control objective is to stabilize the equilibrium located at the origin. While a proportional error bound is highly beneficial close to the origin, it may become rather conservative if the distance to the desired set point, i.e., the equilibrium located at the origin, grows. Then, using uniform bounds on the approximation error, as imposed in the following assumption, may result in tighter bounds. Moreover, we levarage this bound in the following to define admissibility of control sequences.
Assumption 2.
Consider a set $S\subseteq\mathbb{R}^{n}$. Let, for some $\eta>0$, the uniform error bound
$\|f(x,u)-f^{\varepsilon}(x,u)\|\leq\eta$ (5)
hold for all $x\in S$, $u\in\mathbb{U}$, and $\varepsilon\in(0,\bar{\varepsilon}]$ for the approximation of system (1) by the surrogate model (3).
@@
In addition, we require uniform Lipschitz continuity of the surrogate model in the approximation accuracy $\varepsilon$ to derive asymptotic stability of the origin w.r.t. the MPC closed loop.
@@
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior. Let the surrogate dynamics (3) be Lipschitz continuous in the first argument on the set $S$ uniformly in $u\in\mathbb{U}$ and $\varepsilon\in(0,\bar{\varepsilon}]$, i.e., there exists $\bar{L}=\bar{L}(S)$ such that, for every $\varepsilon\in(0,\bar{\varepsilon}]$, there is a Lipschitz constant $L_{f^{\varepsilon}}$ with $L_{f^{\varepsilon}}\leq\bar{L}$ satisfying
$\|f^{\varepsilon}(x,u)-f^{\varepsilon}(y,u)\|\leq L_{f^{\varepsilon}}\|x-y\|$
for all $x,y\in S$ and $u\in\mathbb{U}$.
In Section 4, we verify Assumptions 1, 2, and 3 for data-driven surrogate models (3) generated by kernel EDMD using Koopman theory.
@@
with symmetric and positive-definite weighting matrices $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$. Moreover, we require the following notion of admissibility, which is based on the uniform approximation error of Assumption 2.
Definition 1 (Admissible control sequences on a set $S$).
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior. A control sequence $\mathbf{u}=(u(k))_{k=0}^{N-1}\subset\mathbb{U}$ of length $N$ is said to be admissible for state $\hat{x}\in S$, if $x_{\mathbf{u}}(k)\in S$ holds for all $k\in[1:N]$, where
$x_{\mathbf{u}}(k)=f(x_{\mathbf{u}}(k-1),u(k-1))$
using the dynamics (1) and $x_{\mathbf{u}}(0)=\hat{x}$. For $\hat{x}\in S$, the set of admissible control sequences is denoted by $\mathcal{U}_{N}(\hat{x})$. If, for $\mathbf{u}=(u(k))_{k=0}^{\infty}$, $(u(k))^{N-1}_{k=0}\in\mathcal{U}_{N}(\hat{x})$ holds for all $N\in\mathbb{N}$, we write $\mathbf{u}\in\mathcal{U}_{\infty}(\hat{x})$.
If the surrogate model (3) is considered, we change Definition 1 by tighening the state constraint $x_{\mathbf{u}}(k)\in S$. To be more precise, we impose the set inclusion
$x_{\mathbf{u}}^{\varepsilon}(k)=f^{\varepsilon}(x_{\mathbf{u}}^{\varepsilon}(k-1),u(k-1))\in S\ominus\mathcal{B}_{k\eta}(0),$ (7)
where $\ominus$ denotes the usual Pontryagin set difference. We call these control sequences admissible for $\hat{x}\in S$ w.r.t. the surrogate (3), which will be denoted by $\mathbf{u}\in\mathcal{U}_{N}^{\varepsilon}(\hat{x})$. The constraint tightening (7) ensures that if admissible controls $\mathcal{U}_{N}^{\varepsilon}(\hat{x})$ are applied at the plant, which is governed by the original dynamics (1), then the states are contained in the set $S$ if only Assumption 2 holds.
@@
Algorithm 1 Data-driven MPC on a set $S$
 Input: Horizon $N\in\mathbb{N}$, surrogate $f^{\varepsilon}$, input constraints $\mathbb{U}$, set $S$.  Initialization: Set $k=0$.
@@
$\begin{split}\min_{\mathbf{u}\in\mathcal{U}_{N}^{\varepsilon}(\hat{x})}\quad&\sum\nolimits_{i=0}^{N-1}\ell(x_{\mathbf{u}}^{\varepsilon}(i),u(i))\\ \text{s.t.}\quad&x_{\mathbf{u}}^{\varepsilon}(0)=\hat{x}\text{ and, for $i\in[0:N-1]$,}\\ &x_{\mathbf{u}}^{\varepsilon}(i+1)=f^{\varepsilon}(x_{\mathbf{u}}^{\varepsilon}(i),u(i))\end{split}$ (OCP)
@@
(3) Apply the MPC feedback law $\mu_{N}^{\varepsilon}(\hat{x})=u^{\star}({0})$ at the
      plant, shift $k=k+1$, and go to Step (1).
The (optimal) value function $V^{\varepsilon}_{N}:S\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\}$ for the MPC algorithm associated to the optimal control problem (OCP) is defined by
$V^{\varepsilon}_{N}(\hat{x}):=\inf_{\mathbf{u}\in\mathcal{U}^{\varepsilon}_{N}(\hat{x})}J^{\varepsilon}_{N}(\hat{x},\mathbf{u})$
with $J^{\varepsilon}_{N}(\hat{x},\mathbf{u}):=\sum_{k=0}^{N-1}\ell(x^{\varepsilon}_{\mathbf{u}}(k),u(k))$ and $x^{\varepsilon}_{\mathbf{u}}(k)$ defined in Step (2) of Algorithm 1. Analogously, we define the nominal cost $J_{N}(\hat{x},\mathbf{u}):=\sum_{k=0}^{N-1}\ell(x_{\mathbf{u}}(k),u(k))$, where $x_{\mathbf{u}}(k)$, $k\in[0:N-1]$ is obtained by propagating the original system dynamics (1) starting from $x_{\mathbf{u}}(0)=\hat{x}$. Moreover, we define the nominal optimal value function by
$V_{N}(\hat{x}):=\inf_{\mathbf{u}\in\mathcal{U}^{\varepsilon}_{N}(\hat{x})}J_{N}(\hat{x},\mathbf{u}),$
where we employed the tightened admissible set as well.

### v2

@@ Abstract.
In this paper, we analyze stability of nonlinear model predictive control (MPC) using data-driven surrogate models in the optimization step. First, we establish asymptotic stability of the origin, a controlled steady state, w.r.t. the MPC closed loop without stabilizing terminal conditions for sufficiently long prediction horizons. To this end, we prove that cost controllability of the original system is preserved if sufficiently accurate proportional bounds on the approximation error hold. Here, proportional refers to state and control. The proportionality of the error bounds is a key element to derive asymptotic stability in presence of modeling errors and not only practical asymptotic stability. Second, we exemplarily verify the imposed assumptions for data-driven surrogates generated with kernel extended dynamic mode decomposition based on Koopman operator theory. Hereby, we do not impose invariance assumptions on finite dictionaries, but rather derive all conditions under non-restrictive conditions. Finally, we demonstrate our findings with numerical simulations.
@@ 1. Introduction
Model predictive control (MPC) is an optimization-based control algorithm, that optimizes the future trajectory of the system on a finite horizon and defines a state feedback law by applying to the system the first element of the computed optimal input sequence Rawlings et al. (2017). When, e.g., a first-principles-based description of the system under control is not available, data-driven methods can be used to obtain a surrogate model for the prediction and optimization step. For linear systems, a popular data-driven approach Coulson et al. (2019); Berberich et al. (2020) is based on Willems’ fundamental lemma Willems et al. (2005) avoiding the model identification step and directly using input-output data Markovsky and Dörfler (2021); Faulwasser et al. (2023). For nonlinear systems, data-driven surrogate models can be obtained, e.g., by means of nonlinear ARX models De Nicolao et al. (1997), neural networks Schimperna and Magni (2024); Ren et al. (2022), Bayesian identification Piga et al. (2019) and Gaussian process regression Hewing et al. (2019); Koller et al. (2018).
@@
A numerically and analytically appealing approach is Extended Dynamic Mode Decomposition (EDMD Williams et al. (2015a)) based on the Koopman operator. The Koopman operator is a linear but infinite dimensional operator that encodes the behavior of the associated nonlinear dynamical system Mezić (2005); Rowley et al. (2009). EDMD provides a data-driven approximation of the compression of the Koopman operator to a finite-dimensional subspace spanned by the so-called observable functions, see Mezić (2021) and the recently published overview (Strässer et al., 2026b, Section 2). Various approaches extend this idea to control systems, see, e.g., the recent preprint Haseli et al. (2025) as well as (Strässer et al., 2026b, Section 3). In linear EDMD with control (EDMDc; Proctor et al. (2016); Korda and Mezić (2018)), a surrogate control system, that is linear in the (lifted) state and control, is used. However, as shown in Iacob et al. (2024), this method is often insufficient to capture direct state-control couplings. Bilinear models as in Peitz et al. (2020); Strässer et al. (2026a) offer a reliable alternative that is underpinned by finite-data error bounds Nüske et al. (2023); Schaller et al. (2023). Besides ensuring a sufficient amount of data, a key challenge in EDMD is the choice of the observable functions. Here, data-driven approaches, such as deep neural networks Yeung et al. (2019); Bakarji et al. (2023); Johnson et al. (2025) or kernel EDMD Klus et al. (2020) provide a remedy. In kernel EDMD, the dictionary consists of the canonical features of an a-priori chosen kernel function centered at the data points. For this method, it is possible to build upon the solid approximation-theoretic foundation of reproducing kernel Hilbert spaces (RKHSs) to obtain finite-data error bounds Köhne et al. (2025) leveraging Koopman invariance of suitably chosen RKHSs. This method and its error bounds have been extended to control systems in Bold et al. (2025b). An alternative approach, that shares the advantage of flexible state-control sampling, was proposed in Bevanda et al. (2024) using an additional kernel function to express the dependency on the control input. So far, however, no error bounds are available for this method. Koopman operator-based methods are not the only class of surrogate models for which error bounds can be provided. Other appealing identification technique, that could be used to derive data-driven surrogate models with error bounds include kernel and Gaussian-process regression, see Care et al. (2023); Bisiacco and Pillonetto (2024) and Lederer et al. (2019); Scampicchio et al. (2025), respectively.
@@
The use of data-driven models gives rise to approximation errors, that are related to the choice of the model class used for approximation and the finite amount of data available for the identification. While asymptotic stabilization of a steady-state equilibrium can be guaranteed in the nominal case, i.e., when the model used in the optimization step coincides with the system under control, in presence of modeling errors typically only practical asymptotic stability is achieved, see, e.g., (Grüne and Pannek, 2016, Section 11.5) and Kuntz and Rawlings (2025). Practical asymptotic stability means that the tracking error decreases until a certain threshold and, then, stagnates. Typically, the stagnation can be related proportionally to the approximation error. Practical asymptotic stability of MPC with surrogate models based on EDMD has been shown both without and with the use of terminal ingredients in Bold et al. (2025a) and Worthmann et al. (2024), respectively. A method to obtain offset-free tracking in presence of modeling errors and asymptotically constant disturbances is so-called offset-free MPC Pannocchia and Rawlings (2003); Morari and Maeder (2012). This technique has been successfully within the Koopman framework. In particular, Son et al. (2022) and Chen et al. (2022) consider linear offset-free MPC with surrogate models based on EDMDc, while Schimperna et al. (2025b) considers bilinear EDMD models and addresses the case in which a full description of the reference steady state is not available. Finally, asymptotic stability of the MPC closed-loop can be proven under suitable conditions on the modeling error. In particular, Santos et al. (2008) proves asymptotic stability of an MPC closed-loop with soft constraints and terminal conditions for sufficiently small model-error bounds under the assumption that the terminal constraint is met by the MPC using the model. Similarly, the recent work Kuntz and Rawlings (2025) derives asymptotic stability of MPC with terminal conditions and without state constraints in presence of a sufficiently small parametric error in the model such that the equilibrium in the origin is preserved, see also Schimperna et al. (2025a) for the treatment of state constraints and a rigorous verification of all assumptions using Koopman operator theory. However, asymptotic stability of MPC without terminal conditions in the presence of approximation error has not been studied before.
@@
The first contribution of this work is a general framework for stability analysis of MPC with data-driven surrogate models subject to approximation errors. To be more precise, we show that, if a proportional error bound on the surrogate holds, the MPC algorithm designed using the data-driven surrogate model asymptotically stabilizes the system. With proportional we mean that the error vanishes at the desired set point and its magnitude is bounded by a quantity depending on the size of state and input. We show that proportional error bounds imply the preservation of cost controllability Grimm et al. (2005); Grüne et al. (2010); Coron et al. (2020) in data-driven approximations provided a sufficiently high approximation accuracy. Then, invoking this controllability property, we rigorously show asymptotic stability of the origin w.r.t. the MPC closed loop using the surrogate model in the optimization step if the prediction horizon is sufficiently long, where the origin serves as a prototypical controlled steady state.
@@
The second main contribution of this paper is to verify the assumptions on the data-driven surrogate leveraging recently proposed finite-data error bounds on kernel EDMD under non-restrictive assumptions. To this end, we consider the learning framework proposed in Bold et al. (2025b), suitably modified to ensure that the model is consistent with the system dynamics in the origin. We derive proportional and uniform error bounds – without requiring invariance assumptions on finite dictionaries using a flexible sampling technique to alleviate data requirements. Finally, our findings are verified with numerical simulations. [^1] We point out that this part is an extension of our accepted conference paper Bold et al. (2025c) now properly embedded into a more general learning framework such that we were able to infer asymptotic stability (instead of only practical asymptotic stability). Moreover, we point out that the stability framework for data-driven MPC is independent of the Koopman operator such that it can also be applied with different learning-based techniques.
@@
Notation. The notation $\|\cdot\|$ is used for the Euclidean norm on $\mathbb{R}^{n}$ and its induced matrix norm on $\mathbb{R}^{n\times n}$, while $\|\cdot\|_{F}$ is used to denote the Frobenius norm. Given two numbers $a,b\in\mathbb{N}$, the abbreviation $a:b]:=\mathbb{Z}\cap[a,b]$ is used. Given a set $\Omega$, we denote by $\overline{\Omega}$ its closure and by $\operatorname{int}(\Omega)$ its interior. Given $x_{0}\in\mathbb{R}^{n}$ and $r\in\mathbb{R}_{\geq 0}$, $\mathcal{B}_{r}(x_{0}):=\{x\in\mathbb{R}^{n}:\|x-x_{0}\|\leq r\}$ is the closed ball of radius $r$ centered in $x_{0}$. $\mathcal{C}_{b}(\Omega)\coloneqq\mathcal{C}_{b}(\Omega;\mathbb{R}^{n})$ denotes the space of continuous and bounded functions on the set $\Omega$. The notation $J_{f}$ is used to denote the Jacobian of the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, and $J_{f,x}$ is used to denote the Jacobian of $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, $x\mapsto f(x,u)$ w.r.t. the argument $x$. With $\lambda_{\max}(Q)$ and $\lambda_{\min}(Q)$ we denote the maximum and minimum eigenvalue of the positive definite matrix $Q$, respectively. The symbols $\oplus$ and $\ominus$ represent the Pontryagin set sum and difference, respectively.
@@ 2. Data-driven MPC formulation
with state $x\in\mathbb{R}^{n}$ and input $u\in\mathbb{U}$, where $\mathbb{U}\subseteq\mathbb{R}^{m}$ is a convex and compact set containing the origin, i.e., $0\in\mathbb{R}^{m}$, in its interior. The system dynamics $f$ is assumed to be continuous and locally Lipschitz continuous in its first argument, i.e., for each compact set $K\subset\mathbb{R}^{n}$ there is a Lipschitz constant $L_{f}=L_{f}(K)$ such that, for all $x,y\in K$,
$\|f(x,u)-f(y,u)\|\leq L_{f}\|x-y\|\qquad\forall\,u\in\mathbb{U}.$ (2)
@@
Let the origin be a controlled equilibrium of system (1) for control input $u=0$, i.e. $f(0,0)=0$. The objective is to steer the system to the origin using an MPC controller while taking input constraints $u\in\mathbb{U}$ into account. [^2] The upcoming analysis can be directly applied to the regulation of arbitrary controlled equilibria by considering the shifted dynamics. However, since the system dynamics (1) is assumed to be unknown, the MPC design is based on data-driven surrogate models
$x^{+}=f^{\varepsilon}(x,u)$ (3)
@@
with $\varepsilon\in(0,\bar{\varepsilon}]$ for some $\bar{\varepsilon}>0$. The superscript $\varepsilon$ refers to the approximation accuracy and, throughout the paper, will be used to refer to all the parameters related to the surrogate dynamics (3). Typically, the approximation error depends on the available data (quantity, distribution etc.).
We consider models accompanied by proportional error bounds, where proportional means that the error is zero at the origin and grows, at most, proportionally in the size of state and input. This property is formalized in the following assumption.
Assumption 1 (Error bounds on a set $\Omega$).
Consider a set $\Omega\subseteq\mathbb{R}^{n}$ containing the origin, i.e., $0\in\mathbb{R}^{n}$, in its interior. Let $f^{\varepsilon}$ defined by (3) be a surrogate model for the control system (1). For every $\varepsilon\in(0,\bar{\varepsilon}]$, let the surrogate model satisfy the
  1. (P)
    proportional error bound
    $\|f(x,u)-f^{\varepsilon}(x,u)\|\leq c_{x}^{\varepsilon}\|x\|+c_{u}^{\varepsilon}\|u\|$ (P-bound)
  2. (U)
    uniform error bound
    $\|f(x,u)-f^{\varepsilon}(x,u)\|\leq\eta^{\varepsilon}$ (U-bound)
@@
for all $x\in\Omega$, $u\in\mathbb{U}$ such that the parameters $c_{x}^{\varepsilon}$, $c_{u}^{\varepsilon}$, and $\eta^{\varepsilon}$ satisfy $\lim_{\varepsilon\searrow 0}\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon},\eta^{\varepsilon}\}=0$.
@@
The proportional bound (P-bound) of Assumption 1 requires that the surrogate model is exact at the origin, which is a reasonable assumption since the control objective is to stabilize the equilibrium located at the origin. While a proportional error bound is highly beneficial close to the origin, it may become rather conservative if the distance to the desired set point, i.e., the equilibrium located at the origin, grows. Then, using the uniform bounds (U-bound) of Assumption 1 on the approximation error may result in tighter bounds.
@@
In addition to Assumption 1, we require uniform Lipschitz continuity of the surrogate model in the approximation accuracy $\varepsilon$ to derive asymptotic stability of the origin w.r.t. the MPC closed loop.
@@
Consider a set $\Omega\subseteq\mathbb{R}^{n}$ containing the origin in its interior. Let the surrogate dynamics (3) be locally Lipschitz continuous in the first argument on $\Omega$ uniformly in $u\in\mathbb{U}$ and $\varepsilon\in(0,\bar{\varepsilon}]$, i.e., for every compact set $K\subseteq\Omega$, there exists $\bar{L}=\bar{L}(K)$ such that, for every $\varepsilon\in(0,\bar{\varepsilon}]$, there is a Lipschitz constant $L_{f^{\varepsilon}}$ with $L_{f^{\varepsilon}}\leq\bar{L}$ satisfying
$\|f^{\varepsilon}(x,u)-f^{\varepsilon}(y,u)\|\leq L_{f^{\varepsilon}}\|x-y\|\quad\forall\,x,y\in K,u\in\mathbb{U}.$
In Section 4, we verify Assumptions 1 and 2 for data-driven surrogate models (3) generated by kernel EDMD using Koopman operator theory under suitable smoothness assumptions on the system under control on bounded sets $\Omega$.
@@
with symmetric and positive-definite weighting matrices $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$. Moreover, we require the following notion of admissibility.
Definition 1 (Admissible control sequences).
A control sequence $\mathbf{u}=(u(k))_{k=0}^{N-1}\subset\mathbb{U}$ of length $N\in\mathbb{N}\cup\{+\infty\}$ is said to be admissible. The set of admissible control sequences is denoted by $\mathcal{U}_{N}$.
Admissibility according to Definition 1 does not depend on the system dynamics meaning that the definition is the same for the original dynamics (1) and the surrogate model (3).
@@
Algorithm 1 Data-driven MPC
Input: Horizon $N\in\mathbb{N}$, surrogate $f^{\varepsilon}$, stage cost $\ell$, input constraints $\mathbb{U}$.
Initialization: Set $k=0$.
@@
$\begin{split}\min_{\mathbf{u}\in\mathcal{U}_{N}}(\hat{x})\quad&\sum\nolimits_{i=0}^{N-1}\ell(x_{\mathbf{u}}^{\varepsilon}(i),u(i))\\ \text{s.t.}\quad&x_{\mathbf{u}}^{\varepsilon}(0)=\hat{x}\text{ and, for $i\in[0:N-1]$,}\\ &x_{\mathbf{u}}^{\varepsilon}(i+1)=f^{\varepsilon}(x_{\mathbf{u}}^{\varepsilon}(i),u(i))\end{split}$ (OCP)
@@
(3) Apply the MPC feedback law $\mu_{N}^{\varepsilon}(\hat{x})=u^{\star}({0})$ at the plant to generate the closed loop
$x_{\mu_{N}^{\varepsilon}}(k+1)=f(x_{\mu_{N}^{\varepsilon}}(k),\mu_{N}^{\varepsilon}(x_{\mu_{N}^{\varepsilon}}(k)))$
and shift $k=k+1$, and go to Step (1).
In Algorithm 1, $x_{\mu_{N}^{\varepsilon}}(k)$ denotes the state of the original system governed by the dynamics (1) at time $k$ under the MPC feedback law $\mu_{N}^{\varepsilon}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ computed by solving (OCP) using the data-driven surrogate dynamics $f^{\varepsilon}$.
The (optimal) value function $V^{\varepsilon}_{N}:\Omega\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\}$ associated to the optimal control problem (OCP) is defined by
$V^{\varepsilon}_{N}(\hat{x}):=\inf_{\mathbf{u}\in\mathcal{U}_{N}}J^{\varepsilon}_{N}(\hat{x},\mathbf{u})$
with $J^{\varepsilon}_{N}(\hat{x},\mathbf{u}):=\sum_{k=0}^{N-1}\ell(x^{\varepsilon}_{\mathbf{u}}(k),u(k))$ and $x^{\varepsilon}_{\mathbf{u}}(k)$ defined in Step (2) of Algorithm 1. Analogously, we define the nominal cost $J_{N}(\hat{x},\mathbf{u}):=\sum_{k=0}^{N-1}\ell(x_{\mathbf{u}}(k),u(k))$, where $x_{\mathbf{u}}(k)$, $k\in[0:N-1]$, is obtained by propagating the original system dynamics (1) starting from $x_{\mathbf{u}}(0)=\hat{x}$. The nominal value function is defined as $V_{N}(\hat{x}):=\inf_{\mathbf{u}\in\mathcal{U}_{N}}J_{N}(\hat{x},\mathbf{u})$.